OpenStudy (anonymous):

the observation deck of a skycraper is 420 m above the street. determine the time required for a penny to free-fall from the deck to the street below.

OpenStudy (anonymous):

Time is\[\sqrt{2h/g}\] http://en.wikipedia.org/wiki/Free_fall 9,25 s

OpenStudy (anonymous):

s=ut+1/2at^2, u=0 => s=1/2at^2 a=9.8 s=420 therefore 420/(1/2*9.8) =t^2 => t=9.25s